OpenStudy (anonymous):

12x-96 --------- x^2-4x-32 ---------- 6x^2-6x-12 ------------ 3x+3

OpenStudy (anonymous):

12x-96 12x-96 3x+3 12(x - 8) 3(x + 1) --------- = ------------ * ------------ = -------------- * -------------- x^2-4x-32 x^2-4x-32 6x^2-6x-12 (x + 4)(x - 8) 6(x + 1)(x - 2) ---------- 6x^2-6x-12 ------------ 3x+3 12(x - 8) 3(x + 1) 6 1 6 6 ------------ * -------------- = ------- * ------ = ------------ = -------------- (x + 4)(x - 8) 6(x + 1)(x - 2) x + 4 x - 2 (x + 4)(x - 2) x² + 2x - 8 Factorisation used in solution above x^2-4x-32 = x² -8x + 4x - 32 = x(x - 8) +4(x - 8) = (x + 4)(x - 8) 6x^2-6x-12 = 6(x² - x - 2) = 6[x² -2x + x - 2] = 6[x(x - 2) +1(x - 2)] = 6(x + 1)(x - 2)

OpenStudy (anonymous):

i dont get wwhat the final answer is

OpenStudy (anonymous):

The final answer is 6 ------------ x² + 2x - 8